OpenStudy (anonymous):

Write the equation of the cone z=[square root of:]3x^2+3y^2 in polar coordinates. I don't understand how you convert it.

OpenStudy (anonymous):

Remember that in polar you have \[r^2 = x^2 + y^2\] So \[z = \sqrt{3x^2+3y^2}\rightarrow z = \sqrt{3(x^2 + y^2)} = r\sqrt{3} \]